Please help me solve the following motion problem.
In the Mississippi River near New Orleans, the Creole Queen paddleboat travels 6 miles upstream (against the current) in the same amount of time it travels 12 miles downstream (with the current). If the current of the river is 3 miles per hour, determine the speed of the Creole Queen in still water.
I came up with 1 mile per hour. I checked and it is not the correct answer. I made a table for RT=D but I did not understand where to plug in the figures given.

Answer by checkley77(12844)

D=RT OR T=D/R
T=6/(R-3)
T=12/(R+3)
6/(R-3)=12/(R+3)
12(R-3)=6(R+3)
12R-36=6R+18
12R-6R=18+36
6R=54
R=54/6
R=9 MPH. IS THE SPEED OF THE BOAT IN STILL WATER.
PROOF:
6/(9-3)=12/(9+3)
6/6=12/12
1=1